OpenStudy (mikezack123):

132 = 1/2 x + 2(x - 9) Solve for x

OpenStudy (whpalmer4):

This could have stayed in the original problem, no? \[132 = \frac{1}{2}x + 2(x-9)\]Why don't you multiply both sides by 2, and distribute the expression in the parentheses?

OpenStudy (mikezack123):

whats a parentheses?

OpenStudy (whpalmer4):

( )

OpenStudy (mikezack123):

the original problem says ..... The regular price of a child's entry ticket to a water park is $9 less than that for an adult's. The park offers half off of all entry tickets during the off-peak season. The Sandlers paid a total of $132 for 1 adult ticket and 4 child's tickets to the water park during the off-peak season. The following equation represents this situation, where x represents the regular price of an adult ticket : 132 = 1/2 x + 2(x - 9)

OpenStudy (mikezack123):

What is the regular price of a child's ticket? $60 $57 $51 $42

OpenStudy (whpalmer4):

one parenthesis, two parentheses right. I talked to you the previous time you posted this, when I spotted you'd made a mistake copying the equation... So, again, multiply the entire equation by 2. This will eliminate the fraction. Then, use the distributive property of multiplication to get rid of the stuff in the parentheses. Show me what you get.

OpenStudy (anonymous):

x=60

OpenStudy (mikezack123):

i dont get it.

OpenStudy (whpalmer4):

Okay, I'll spell it out: \[132 = \frac{1}{2}x+2(x-9)\]Multiply by 2 \[2*132 = 2*\frac{1}{2}x+2*2(x-9)\]\[264 = x+4(x-9)\]Now distribute the right hand side \[264=x+4x-4*9\]\[264=5x-36\]Now add 36 to each side\[264+36=5x-36+36\]\[300=5x\]Can you solve that?